OpenStudy (anonymous):

If a sphere has a radius of 4 inches, what is its surface area? volume? (Round your answers to two decimal places.)

OpenStudy (anonymous):

surface area of a sphere = 4 x pi x radius squared 4 x pi x 4squared = 201.06 inches

OpenStudy (anonymous):

volume of a sphere = 4/3 x pi x radius cubed = 4/3 x pi x 4 cubed = 268.08

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

dont mention it